Roger can hang wallpaper on a square wall with 12-ft sides in 40 minutes. Assuming the wallpaper is hung at the same rate, how long will it take him to cover a square section of wall with 6-ft sides?


Thanks!

Answer by checkley71(8403)

12^2/40=6^2/X OR 144/40=36/X OR 144X=1440 OR X=1440/144 OR X=10 MINUTES.